Find the centroid of the region.



The triangle with vertices (0, 0), (2, 0), and (0, 1).


1
Expert's answer
2022-05-03T07:30:05-0400

"x_{0}=\frac{x_{1}+x_{2}+x_{3}}{3}, y_{0}=\frac{y_{1}+y_{2}+y_{3}}{3}"

"x_0 = \frac{0+0+2}{3}=\frac{2}{3}, y_0=\frac{0+0+1}{3}=\frac{1}{3}"